Triangle:
A-B= 65
A-C= 33
B-C= 56
Identify sin A in simplest terms.
Identify cos A in simplest terms.
Identify tan A in simplest terms.
Please help me on this I could not draw the triangle. I could just give the numbers. Thank You

Answer by stanbon(75887)

Triangle:
A-B= 65
A-C= 33
B-C= 56
Identify sin A in simplest terms.
Identify cos A in simplest terms.
Identify tan A in simplest terms.
-------------------------
1st use the Law of cosines to find cos(A)
The side opposite A is B-C
---------
cos(A) = (b^2 + c^2-a^2)/2bc
cos(A) = (33^2 + 65^2 - 56*2)/(2*33*65)
cos(A) = 0.5077
-----------------------
A = cos^-1(0.5077) = 59.5 degrees
---------------------------------------
2nd sin(A) = sin(59.5) = 0.8615
-----------------------
3rd tan(A) = sin(A)/cos(A) = 0.8615/0.5077 = 1.697
=========================================================
Cheers,
Stan H.